# Ceresa cycles of $X_{0}(N)$

### v1

@@ Abstract.
The Ceresa cycle is an algebraic 1-cycle on the Jacobian of an algebraic curve. Although it is homologically trivial, Ceresa famously proved that for a very general complex curve of genus at least 3, it is non-trivial in the Chow group. In this paper we study the Ceresa cycle attached to the complete modular curve $X_{0}(N)$. For prime level $p$ we give a complete description, namely we prove that if $X_{0}(p)$ is not hyperelliptic, then its Ceresa cycle is non-torsion. For general level $N$, we prove that there are finitely many $X_{0}(N)$ with torsion Ceresa cycle. Our method relies on the relationship between the vanishing of the Ceresa cycle and Chow-Heegner points on the Jacobian. We use the geometry and arithmetic of modular Jacobians to prove that such points are of infinite order and therefore deduce non-vanishing of the Ceresa cycle.
@@ 1. Introduction
which is an element of the Chow group $CH_{1}(J)$. This cycle is always homologically trivial, that is, it lies in the kernel of the cycle class map form $CH_{1}(J)$ into a Weil cohomology theory, but crucially it is not always trivial in the Chow group. In fact, for a very general complex curve of genus at least 3, Ceresa [7] proved that it is algebraically, and in particular rationally, non-trivial independent of the choice of base point divisor.
@@
In this paper we study a classically interesting family of curves, namely the modular curves whose non-cuspidal points parametrise isomorphism classes of elliptic curves with a marked isogeny. We study the Ceresa cycle of $X_{0}(N)$ inside $\operatorname{CH}_{1}(J)\otimes\mathbb{Q}$, henceforth simply $\operatorname{CH}_{1}(J)$. As detailed in Theorem 3 (c.f. [39]), the Ceresa cycle can only possibly vanish for one choice of basepoint divisor class, namely $e=\frac{K_{C}}{2g-2}$, where $K_{C}$ is the canonical class and $g$ is the genus of the curve. For the rest of this paper we write $\operatorname{Cer}(N)$ for $\operatorname{Cer}\left(X_{0}(N),\frac{K_{C}}{2g-2}\right)$. Recently, a large family of modular curves was studied by Kerr, Li, Qiu and Yang [23]. This work proves that the Ceresa cycle associated to the complete modular curve corresponding to the congruence subgroup $\Gamma(2)\cap\Gamma_{1}(2N)$ is non-zero in $\operatorname{CH}_{1}(J)$, for $N$ sufficiently large. We remark that their results are implied by our work, combined with a statement about Ceresa cycles and covers (see Proposition 24). The main results of our paper are the following.
@@
then $\operatorname{Cer}(N)$ is non-zero in $CH_{1}(J)$.
We note that Theorem 2 combined with Theorem 3 (reproduced from [39]) proves a significant part of [33, Conjecture 1.2.2].
@@
Our methodology is similar to [23], as we reduce non-vanishing to the study of a certain divisor in the modular Jacobian. However, the key difference is that the congruence subgroup considered in the previous work allows the authors to reduce to Heegner divisors and conclude that such points are of infinite order by using the celebrated Gross-Zagier formula to show that the Néron-Tate height of the divisor is non-zero. In our case, the associated points do not fall into a pre-existing framework, and we instead make use of the arithmetic and geometry of the modular Jacobian to deduce that the associated point is non-torsion. We do not believe such points have received in-depth study in the modular curves literature to date, as despite being linear combinations of Heegner divisors, they do not satisfy the assumptions of the Gross-Zagier height theorem [20].
One drawback of our method (and of [23]) is that it does not apply to examples where $J_{0}(N)(\mathbb{Q})$ has rank 00. A famous examples of this is $X_{0}(64)$, which is simply the Fermat quartic, whose Jacobian has $\mathbb{Q}$-rank zero, but for which the Ceresa cycle does not vanish [21]. There are other examples excluded by bound of Theorem 2 with finite $J_{0}(N)(\mathbb{Q})$, for instance $N=34,38,51,55\ldots$. The curve $X_{0}(34)$ has genus $3$ and we were able to verify non-vanishing using [16], however we were unable to apply this to higher genus curves due to the automorphism hypothesis. Moreover, our examples fail the cohomological vanishing criteria proved in [26].
@@
Beyond constructing new points of infinite order on the modular Jacobian, the main arithmetic application of our work is towards triple product $L$-functions. Namely, there are relations between the height of $(f,g,h)$-isotypic Hecke components of the Gross-Kudla-Schoen cycle (an object closely related to the Ceresa cycle e.g. [39]), and the special value $L^{\prime}(f\times g\times h,2)$, by analogy with the Gross-Zagier height formula [20] [38] [2]. Moreover, similar points are also studied by Dogra and Le Fourn [14] in the context of the quadratic Chabauty method for modular curves.
This paper is organised as follows. In Section 2, we recall some standard constructions associated to Chow groups and define ‘shadow points’ corresponding to Hecke operators on $J_{0}(N)$. In Section 3, we give a brief overview of some well-known results on the geometry and arithmetic of modular curves which allow us to study the shadow points. In Sections 4 and 5, we prove Theorems 1 and 2 respectively.
@@ Acknowledgements.
The authors thank Vladimir Dokchitser, Jef Laga, Samir Siksek and Jan Vonk for helpful conversations. Both authors are grateful for the financial support provided by the UK Engineering and Physical Sciences Research Council. The first named author is supported by the EPSRC Doctoral Prize fellowship EP/W524335/1. The second named author is supported by the EPSRC studentship EP/W523793/1.
@@ 2. Gross-Kudla-Schoen Cycles and Chow-Heegner Points
In what follows $C$ and $D$ are smooth, projective and geometrically integral curves, defined over an algebraically closed field $K$. We write $J_{C}$ for the Jacobian variety of $C$, and we denote the genus of $C$ by $g$.
@@ 2.1. The Ceresa Cycle and Gross-Kudla-Schoen Classes
In light of this result, we restrict to the case $e=\frac{K_{C}}{2g-2}$ and write $\operatorname{Cer}(C)$ for $\operatorname{Cer}(C,\frac{K_{C}}{2g-2})$, for any choice of canonical divisor.
@@ 2.2. Chow-Heegner and Shadow Points
Recall that to any morphism of Jacobians, $\phi:J_{C}\to J_{D}$, there is a corresponding correspondence, $C\leftarrow X\rightarrow D$, which realises the isogeny via push-forward and pullback. When $\phi$ is an endomorphism, this correspondence can be naturally embedded in $C^{2}$, and we may refer to fixed points of $\phi$ as the intersection of the correspondence with the diagonal.
@@
We remark that this divisor is a scaling of the corresponding Chow-Heegner point, see [10], [9] for details on the construction of such points. The terminology of “shadows” is adopted from [16], both due to the similarity of the construction, and to avoid any notational ambiguity with the Chow group. If $\phi$ is defined over $\mathbb{Q}$, then so is $\operatorname{Sh}(\phi)$.
The following proposition is similar to those found in [16] and [23].
@@
Suppose $J_{C}$ has a non-trivial endomorphism,$\phi$. If the Ceresa cycle is zero in $\operatorname{CH}_{1}(J)$, then $[\operatorname{Sh}(\phi)]\in J_{C}$ is torsion.
@@
As mentioned before this proposition, associated to $\phi$ there is a natural class $X_{\phi}$ in $C^{2}$, given by the image of the correspondence realising $\phi$. We can push this class forward through $\Gamma^{3}(C,e)$, by the following $\operatorname{pr}_{3,*}\left(\Gamma^{3}(C,e)\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right)$, where $e=\frac{K_{C}}{2g-2}$. This can be calculated by computing this pushforward for each term in the definition of $\Gamma^{3}(C,e)$. We summarise these calculations in Table 1, using the notation of the previous definition.
@@
Combining these terms, we get the following point, $F_{\phi}-\deg(F_{\phi})e-\phi(e)-\phi^{\vee}(e)+\deg(\phi)e+\deg(\phi^{\vee})e$, which is commonly the definition of a Chow-Heegner point. Multiplying by $22g-$ 2 to clear denominators gives $\operatorname{Sh}(\phi)=(2g-2)F-\deg(F)K_{C}-\phi(K_{C})-\phi^{\vee}(K_{C})+(\deg(\phi)+\deg(\phi^{\vee}))K_{C}$. If the Ceresa cycle is zero in the rational Chow group, then so is $\Gamma^{3}(C,e)$, and so pushing forward $X_{\phi}$ through $\Gamma^{3}(C,e)$ must also be zero. This shows the resulting divisor is trivial in the Jacobian tensored with $\mathbb{Q}$, and so it is torsion. ∎
2.3. Fixed Points of Hecke Operators
In the case of modular curves there are many non-trivial endomorphisms of the Jacobian, namely Hecke operators. More precisely, for $C=X_{0}(N)$ the endomorphism ring of the Jacobian contains the Hecke algebra generated (over $\mathbb{Z}$) by the induced Atkin-Lehener operators $w_{d}$ for all $d|N$ and coprime to $\frac{N}{d}$; and for all $\ell\nmid N$ the Hecke operator $T_{\ell}$.
Notation.
(1) $\operatorname{Sh}(\phi)=(2g-2)F_{\phi}-\deg(F)K_{C}.$
The fixed points of Atkin-Lehner operators are well-know due to Ogg [31, Proposition 3].
@@ 2.3. Fixed Points of Hecke Operators
The points in the support of $F_{w_{d}}$ are either:
  • cusps if $d=4$;
  • CM points with endomorphism ring containing $\mathbb{Z}[\sqrt{-d}]$
Moreover, all fixed points occur with multiplicity $1$.
${X_{0}(N\ell)}$${X_{0}(N\ell)}$${X_{0}(N)}$${X_{0}(N)}$$\scriptstyle{w_{\ell}}$
see [13, Chapter 5] for details. As a result of this, we see that in the sense of the previous section, $T_{\ell}=T_{\ell}^{\vee}$. Further, as is well known, the degrees for both are $\ell+1$. This simplifies $\operatorname{Sh}(T_{\ell})$ to $(2g-2)F_{T_{\ell}}-\deg(F_{T_{\ell}})K_{X_{0}(N)}-2T_{\ell}(K_{X_{0}(N)})+2(\ell+1)K_{X_{0}(N)}$.
Proposition 6.
The points in the support of $F_{T_{\ell}}$ are either:
  • Cusps;
@@
Proposition 7.
Suppose $P\in X_{0}(N\ell)$ has both projections to $X_{0}(N)$ equal, then the image of the tangent space of $P$ under the map from $X_{0}(N\ell)$ to $X_{0}(N)^{2}$ is not diagonal, for $\ell=2,3,5$.
Proof.
Otherwise, $P$ corresponds to an elliptic curve. By the previous proposition, this elliptic curve has CM. Since the $j$-invariant is a local parameter for CM points on $X_{0}(N)$, it is enough to check that $j-j\circ\omega_{\ell}$ has simple zeros at CM points on $X_{0}(N\ell)$. As this function is a pullback from $X_{0}(\ell)$, we start by proving this for $X_{0}(\ell)$. For $\ell=2$, $j-j\circ\omega_{\ell}$ has two poles of order 2, and so is a degree 4 function. There are 3 CM orders with an element of norm 2, $\mathbb{Z}[\sqrt{-1}]$, $\mathbb{Z}[\sqrt{-2}]$ and $\mathbb{Z}[\frac{-1+\sqrt{-7}}{2}]$. The first two of these have a unique ideal of index 2, and the second has 2 choices, giving rise to a total of 4 points on $X_{0}(2)$ where this function vanishes. This forces all 4 points to be simple zeros. Similar counting arguments hold for $\ell=3,5$, where for $\ell=5$, class numbers also have to be taken into account. It remains to show that the simple zeros on $X_{0}(\ell)$ pullback to simple zeros on $X_{0}(N\ell)$. This can only fail if $X_{0}(N\ell)\to X_{0}(\ell)$ ramifies over these points, but this does not happen (e.g. by considering the order of vanishing of $j$). ∎
This result should hold more generally by considering fist order deformations of lattices, however, as we only use these 3 cases, we simply prove it for these, mainly for exposition reasons.
Proposition 8.
Proof.
Combining the above in the special case of $T_{2}$, gives the following
$F_{T_{2}}=D_{4}+D_{8}+2D_{7}+2D_{\infty}$

### v2

@@ Abstract.
The Ceresa cycle is an algebraic 1-cycle on the Jacobian of an algebraic curve. Although it is homologically trivial, Ceresa famously proved that for a very general complex curve of genus at least 3, it is non-trivial in the Chow group. In this paper we study the Ceresa cycle attached to the complete modular curve $X_{0}(N)$ modulo rational equivalence. For prime level $p$ we give a complete description, namely we prove that if $X_{0}(p)$ is not hyperelliptic, then its Ceresa cycle is non-torsion. For general level $N$, we prove that there are finitely many $X_{0}(N)$ with torsion Ceresa cycle. Our method relies on the relationship between the vanishing of the Ceresa cycle and Chow-Heegner points on the Jacobian. We use the geometry and arithmetic of modular Jacobians to prove that such points are of infinite order and therefore deduce non-vanishing of the Ceresa cycle.
@@ 1. Introduction
which is an element of the Chow group $CH_{1}(J)$, modulo rational equivalence. This cycle is always homologically trivial, that is, it lies in the kernel of the cycle class map form $CH_{1}(J)$ into a Weil cohomology theory, but crucially it is not always trivial in the Chow group. In fact, for a very general complex curve of genus at least 3, Ceresa [9] proved that it is algebraically, and in particular rationally, non-trivial independent of the choice of base point divisor.
@@
In this paper we study a classically interesting family of curves, namely the modular curves whose non-cuspidal points parametrise isomorphism classes of elliptic curves with a marked isogeny. We study the Ceresa cycle of $X_{0}(N)$ inside $\operatorname{CH}_{1}(J_{0}(N))\otimes\mathbb{Q}$, henceforth simply $\operatorname{CH}_{1}(J_{0}(N))$. As detailed in Theorem 3 (c.f. [44]), the Ceresa cycle can only possibly vanish for one choice of basepoint divisor class, namely $e=\frac{K_{C}}{2g-2}$, where $K_{C}$ is the canonical class and $g$ is the genus of the curve. For the rest of this paper we write $\operatorname{Cer}(N)$ for $\operatorname{Cer}\left(X_{0}(N),\frac{K_{C}}{2g-2}\right)$. Recently, a large family of modular curves was studied by Kerr, Li, Qiu and Yang [27]. That work proves that the Ceresa cycle associated to the complete modular curve corresponding to the congruence subgroup $\Gamma_{N}:=\Gamma(2)\cap\Gamma_{1}(2N)$ is non-zero in $\operatorname{CH}_{1}(J_{\Gamma_{N}})$, for $N$ sufficiently large. We remark that their results are implied by our work, combined with a statement about Ceresa cycles and covers (see Proposition 25). The main results of our paper are the following.
@@
then $\operatorname{Cer}(N)$ is non-zero in $CH_{1}(J_{0}(N))$.
We note that Theorem 2 combined with Theorem 3 (reproduced from [44]) proves a significant part of [38, Conjecture 1.2.2].
@@
Our methodology is similar to [27], as we reduce non-vanishing to the study of a certain divisor in the modular Jacobian. However, the key difference is that the congruence subgroup considered in the previous work allows the authors to reduce to Heegner divisors and conclude that such points are of infinite order by using the celebrated Gross-Zagier formula to show that the Néron-Tate height of the divisor is non-zero. In our case, the associated points do not fall into a pre-existing framework, and we instead make use of the arithmetic and geometry of the modular Jacobian to deduce that the associated point is non-torsion.
We remark that Chow-Heegner divisor were also studied by Dogra and Le Fourn [16] in the context of the quadratic Chabauty method for modular curves. In this work, the authors were able to show that certain linear combinations of Chow-Heegner points are torsion using the Gross-Zagier formula and its generalisation.
One drawback of our method (and of [27]) is that it does not apply to examples where $J_{0}(N)(\mathbb{Q})$ has rank 00. A famous examples of this is $X_{0}(64)$, which is simply the Fermat quartic, whose Jacobian has $\mathbb{Q}$-rank zero, but for which the Ceresa cycle does not vanish [25]. There are other examples excluded by bound of Theorem 2 with finite $J_{0}(N)(\mathbb{Q})$, for instance $N=34,38,51,55\ldots$. The curve $X_{0}(34)$ has genus $3$ and we were able to verify non-vanishing using [18], however we were unable to apply this to higher genus curves due to the automorphism hypothesis. Moreover, our examples fail the cohomological algebraic vanishing criteria proved in [30].
@@
Beyond constructing new points of infinite order on the modular Jacobian, the main arithmetic application of our work is towards triple product $L$-functions. Namely, there are relations between the height of $(f,g,h)$-isotypic Hecke components of the Gross-Kudla-Schoen cycle (an object closely related to the Ceresa cycle e.g. [44]), and the special value $L^{\prime}(f\times g\times h,2)$, by analogy with the Gross-Zagier height formula [24] [43] [3].
This paper is organised as follows. In Section 2, we recall some standard constructions associated to Chow groups and define ‘shadow points’ corresponding to endomorphisms of Jacobians. In Section 3, we give a brief overview of some well-known results on the geometry and arithmetic of modular curves which allow us to study the shadow points. In Section 4 we give explicit formulae for shadow points corresponding to Hecke operators on $J_{0}(N)$. In Sections 5 and 6, we prove Theorems 1 and 2 respectively. In Section 7 we discuss the Hecke structure of shadow points and the links between our work and the Gross-Kudla conjecture.
@@ Acknowledgements.
The authors thank Vladimir Dokchitser, Jef Laga, Samir Siksek and Jan Vonk for helpful conversations; and Benedict Gross for his comments on the previous version of this work which have inspired Section 7. Both authors are grateful for the financial support provided by the UK Engineering and Physical Sciences Research Council. The first named author is supported by the EPSRC Doctoral Prize fellowship EP/W524335/1. The second named author is supported by the EPSRC studentship EP/W523793/1.
@@ 2. Gross-Kudla-Schoen Cycles and Chow-Heegner Points
In what follows $C$ and $D$ are smooth, projective and geometrically integral curves, defined over an algebraically closed field $K$. We write $J_{C}$ for the Jacobian variety of $C$ and we denote the genus of $C$ by $g$.
@@ 2.1. The Ceresa Cycle and Gross-Kudla-Schoen Classes
In light of this result, we restrict to the case $e=\frac{K_{C}}{2g-2}$ and write $\operatorname{Cer}(C)$ for $\operatorname{Cer}(C,\frac{K_{C}}{2g-2})$, for any choice of representative of the canonical class.
@@ 2.2. Chow-Heegner and Shadow Points
To any morphism of Jacobians, $\phi:J_{C}\to J_{D}$, there is a correspondence of curves, $C\leftarrow X\rightarrow D$, which realises the isogeny via push-forward and pullback. When $\phi$ is an endomorphism, this correspondence can be naturally embedded in $C^{2}$, and we may refer to fixed points of $\phi$ as the intersection of the correspondence with the diagonal.
@@
We remark that this divisor is a scaling of the corresponding Chow-Heegner point, see [12], [11] for details on the construction of such points. The terminology of “shadows” is adopted from [18], both due to the similarity of the construction and to avoid any notational ambiguity with the Chow group. If $\phi$ is defined over $\mathbb{Q}$, then so is $\operatorname{Sh}(\phi)$. The following proposition is similar to those found in [18] and [27].
@@
Suppose $J_{C}$ has a non-trivial endomorphism $\phi$. If the Ceresa cycle is zero in $\operatorname{CH}_{1}(J)$, then $[\operatorname{Sh}(\phi)]\in J_{C}$ is torsion.
@@
As mentioned before this proposition, associated to $\phi$ there is a natural class $X_{\phi}$ in $C^{2}$, given by the image of the correspondence realising $\phi$. We can push this class forward through $\Gamma^{3}(C,e)$, in the following $\operatorname{pr}_{3,*}\left(\Gamma^{3}(C,e)\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right)$, where $e=\frac{K_{C}}{2g-2}$. This can be calculated by computing this pushforward for each term in the definition of $\Gamma^{3}(C,e)$. We summarise these calculations in Table 1 using the notation of the previous definition.
@@
Combining these terms, we get the following point
$F_{\phi}-\deg(F_{\phi})e-\phi(e)-\phi^{\vee}(e)+\deg(\phi)e+\deg(\phi^{\vee})e$
which is commonly the definition of a Chow-Heegner point. Multiplying by $22g-$ 2 to clear denominators gives $\operatorname{Sh}(\phi)=(2g-2)F-\deg(F)K_{C}-\phi(K_{C})-\phi^{\vee}(K_{C})+(\deg(\phi)+\deg(\phi^{\vee}))K_{C}$. If the Ceresa cycle is zero in the rational Chow group, then so is $\Gamma^{3}(C,e)$, and so pushing forward $X_{\phi}$ through $\Gamma^{3}(C,e)$ must also be zero. This shows the resulting divisor is trivial in the Jacobian tensored with $\mathbb{Q}$, and so it is torsion. ∎
We remark that the above proposition does apply to $\phi$ being the identity map. In this case, $F_{\phi}$ is $-K_{C}$, and the corresponding shadow is $(2g-2)e-K_{C}$, which is trivial by assumption.
2.3. Action of the Endomorphism Algebra
By assumption, $C$ admits a non-trivial correspondence. It is natural to ask how the action of these on the diagonal cycle affect the shadows. We therefore expand the definition of shadow points as follows.
Definition 3.
Let $Z\in CH_{1}(C^{3})\otimes\mathbb{C}$ and let $\phi$ be a correspondence from $C$ to $C$. Define the shadow $\operatorname{Sh}(Z,\phi)$ of $Z$ under $\phi$ to be
$(2g-2)Z_{*}X_{\phi}\in J_{C}\otimes\mathbb{C}$
where as before, $X_{\phi}$ is the graph of $\phi$ in $C\times C$.
@@ 2.3. Action of the Endomorphism Algebra
Fix a correspondence $\phi$ from $C$ to $C$, and take another (not necessarily distinct) correspondence $\psi$, also from $C$ to $C$. Then the shadow points satisfy the following.
  • $\operatorname{Sh}((\psi,\mathrm{id},\mathrm{id})_{*}Z,\phi)=\operatorname{Sh}(Z,\phi\circ\psi)$
  • $\operatorname{Sh}((\mathrm{id},\psi,\mathrm{id})_{*}Z,\phi)=\operatorname{Sh}(Z,\psi^{\vee}\circ\phi)$
  • $\operatorname{Sh}((\mathrm{id},\mathrm{id},\psi)_{*}Z,\phi)=\psi^{\vee}_{*}\operatorname{Sh}(Z,\phi)$
@@
For the first equality we note that
$((\psi,\mathrm{id},\mathrm{id})_{*}Z)_{*}X_{\phi}=\operatorname{pr}_{3,*}\left((\psi,\mathrm{id},\mathrm{id})_{*}Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right).$
Writing $\psi$ as the composition $\pi_{2,*}\pi_{1}^{*}$, where $\pi_{1}:X\to C$ and $\pi_{2}:X\to C$, we get the following
$((\psi,\mathrm{id},\mathrm{id})_{*}Z)_{*}X_{\phi}=\operatorname{pr}_{3,*}\left((\pi_{2},\mathrm{id},\mathrm{id})_{*}(\pi_{1},\mathrm{id},\mathrm{id})^{*}Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right).$
Applying the projection formula to $(\pi_{2},\mathrm{id},\mathrm{id})_{*}$
$((\psi,\mathrm{id},\mathrm{id})_{*}Z)_{*}X_{\phi}=\operatorname{pr}_{3,*}(\pi_{2},\mathrm{id},\mathrm{id})_{*}\left((\pi_{1},\mathrm{id},\mathrm{id})^{*}Z\cdot(\pi_{2},\mathrm{id},\mathrm{id})^{*}\operatorname{pr}_{1,2}^{*}X_{\phi}\right)$
$=\operatorname{pr}_{3,*}\left((\pi_{1},\mathrm{id},\mathrm{id})^{*}Z\cdot(\pi_{2},\mathrm{id},\mathrm{id})^{*}\operatorname{pr}_{1,2}^{*}X_{\phi}\right).$
By the same process in reverse
$\operatorname{pr}_{3,*}\left((\pi_{1},\mathrm{id},\mathrm{id})^{*}Z\cdot(\pi_{2},\mathrm{id},\mathrm{id})^{*}\operatorname{pr}_{1,2}^{*}X_{\phi}\right)=$
$\operatorname{pr}_{3,*}(\pi_{1},\mathrm{id},\mathrm{id})_{*}\left((\pi_{1},\mathrm{id},\mathrm{id})^{*}Z\cdot(\pi_{2},\mathrm{id},\mathrm{id})^{*}\operatorname{pr}_{1,2}^{*}X_{\phi}\right)=\operatorname{pr}_{3,*}\left(Z\cdot(\psi^{\vee},\mathrm{id},\mathrm{id})_{*}\operatorname{pr}_{1,2}^{*}X_{\phi}\right).$
As the action of $(\psi^{\vee},\mathrm{id},\mathrm{id})$ and $\operatorname{pr}_{1,2}$ form a Cartesian square, this can be simplified to $Z_{*}\left((\psi^{\vee},\mathrm{id})_{*}X_{\phi}\right)$, and this is last class is then $X_{\phi\circ\psi}$. The argument for the second is nearly identical.
For the third equality
$((\mathrm{id},\mathrm{id},\psi)_{*}Z)_{*}X_{\phi})=\operatorname{pr}_{3,*}\left((\mathrm{id},\mathrm{id},\pi_{2})_{*}(\mathrm{id},\mathrm{id},\pi_{2})^{*}Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right).$
Using the projection formula, and noting $\operatorname{pr}_{1,2}\circ(\mathrm{id},\mathrm{id},\pi_{2})=\operatorname{pr}_{1,2}$ gives
$\operatorname{pr}_{3,*}\left((\mathrm{id},\mathrm{id},\pi_{2})_{*}(\mathrm{id},\mathrm{id},\pi_{2})^{*}Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right)=\operatorname{pr}_{3,*}(\mathrm{id},\mathrm{id},\pi_{2})_{*}\left((\mathrm{id},\mathrm{id},\pi_{1})^{*}Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}\right).$
Similarly, this is equal to
$\operatorname{pr}_{3,*}(\mathrm{id},\mathrm{id},\pi_{2})_{*}(\mathrm{id},\mathrm{id},\pi_{2})^{*}(Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi})=\operatorname{pr}_{3,*}(\mathrm{id},\mathrm{id},\psi^{\vee})_{*}(Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi}).$
As $(\mathrm{id},\mathrm{id},\psi^{\vee})$ and $\operatorname{pr}_{3}$ give rise to a Cartesian square, we get this is $\psi^{\vee}_{*}\operatorname{pr}_{3,*}(Z\cdot\operatorname{pr}_{1,2}^{*}X_{\phi})$. ∎
These formulae are especially interesting when $\psi$ belongs to a commutative subalgebra of the full endomorphism ring, and $Z$ is taken as a eigen-component of $\Gamma^{3}(C,e)$ for this sub-algebra, as this will give equalities between certain shadows (see Section 7).